# Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA

### v4

@@ Abstract
Long-form videos that span across wide temporal intervals are highly information redundant and contain multiple distinct events or entities that are often loosely related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature explore the use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector. Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, IntentQA. The code can be found at https://github.com/jongwoopark7978/LVNet
@@ 1 Introduction
Video understanding is a long-standing vision problem [1] with numerous real-world applications. It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos [3, 15, 33, 13, 32]. Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos [34, 51, 44].
@@
The recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures [23, 21, 7], has resulted in significant improvements across visual question answering (VQA) tasks. These models demonstrate exceptional performance within the image domain, and their video variants [47, 28, 27] perform similarly on shorter videos, yet demonstrating limited performance on long-form video benchmarks [24, 16, 30, 31]. This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning. An alternate line of works [52, 36, 16, 37] adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain. However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames. Such a design choice leading to high compute expense, is questioned in [4, 30, 37], and is the key motivation for our exploration of key frame selection, i.e. identifying a minimal set of frames most useful for correctly answering a given video-question pair.
Figure 1: (left) Our proposed LVNet contains a novel Hierarchical Keyframe Selector module to select keyframes and followed by VLM and LLM for caption and answer generation. Hierarchical Keyframe Selector initially begins by processing dense frames and keywords with the lighter module and progressively exploits heavier and more performance-oriented modules on a small set of frames, resulting from the reduction of keyframe candidates, to ensure efficient computation. (right) LVNet achieves state-of-the-art performance on the EgoSchema subset while utilizing only a fraction of captioned frames. In particular, LVNet obtains its highest accuracy of 68.2% with 12 captions, outperforming VideoAgent, a model using a similar-scale captions, by 8%. A more detailed analysis of accuracy vs. the number of captions is provided in 4.3.
@@
Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM and LLM for caption and answer generation as illustrated in Figure 1. Aligned with prior work [52, 39, 37], the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair. The scope of this work focuses on optimizing the prior two stages.
We summarize our key contributions as follows:
  • Novel Hierarchical Keyframe Selector module which are composed of three submodules: Temporal Scene Clustering (TSC), Coarse Keyframe Detector (CKD), and Fine Keyframe Detector (FKD).
    • Temporal Scene Clustering performs non-uniform frame sampling by clustering visually similar frames, minimizing redundancy in long videos and capturing key scenes. The lightweight module is used to efficiently filter dense frames.
    • Coarse Keyframe Detector generates keywords, representing atomic activities, using the given query and LLM. It assigns confidence scores to frames based on keyword relevance, sub-sampling high-confidence frames for improved interpretability over visual-only selection.
    • Fine Keyframe Detector refines frame selection by combining multiple frames using visual templating and Visual Language Model (VLM), enabling higher-level reasoning and natural language output for improved accuracy over the CKD’s keyword-based selection.
@@
  • Long-form Video Understanding framework requiring zero video-level training.
@@
Our proposed LVNet achieves state-of-the-art results compared to the models that utilize a similar number of captions on three long-form video question answering benchmarks—EgoSchema[24], NExT-QA[41], and IntentQA[22] as described in 4.2. This demonstrates the strong performance and general applicability of our approach.
@@ 2 Related Work
Video Question Answering: Visual question answering (VQA), a relatively recent task, involves generating free-form and open-ended textual content conditioned on an image and natural language text [2]. Its video variant, Video-VQA [49] replaces images with videos. While multiple early datasets focus on querying objects or events based on referential and spatial relations [44, 51, 49], later tasks require explicit temporal understanding of sequential events [18, 19, 50]. More recent datasets focus on longer videos containing multiple actions and scenes spread over a wide time interval (termed long-form videos) [41, 20] with questions additionally requiring causal reasoning to correctly answer. Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning [41] over long temporal windows [24]. Some works tackling such casual video VQA tasks leverage graph networks to model cross object / event relations [14, 42, 43]. A more recent line of works integrate large-language models (LLMs) to tackle this task [52, 36, 16, 37, 30, 39, 9] utilizing the strong reasoning skills of the LLMs. A common aspect is the use of a vision langauge model (VLM) to convert the frame level visual information into natural language. This is in turn input to the LLM which makes a final prediction.
Unlike these methods, our LVNet incorporates a unique Hierarchical Keyframe Selector that progressively reduces the number of keyframe candidates. Lighter modules are applied to dense frames, while heavier, more performance-focused modules are applied to a small set of frames for efficient computation. Additionally, LVNet does not require video-level training, unlike earlier supervised approaches.
Frame Selection in Videos: The task of frame selection in videos has been long explored in video [8, 53] with more recent works focused directly on long-form video question answering [4, 30]. Most similar to our work is [37] which employs an LLM based strategy for video frame selection. However, our proposed Hierarchical Keyframe Selector module differs with unique visual only operations as well as cross-modal operations handling much longer frame sequences extending up to 900 frames.
@@ 3 Method
Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Vision Language Model (VLM), and a Large Language Model (LLM) as illustrated in Figure 1. The HKS, an efficient, hierarchical keyframe selector, is the core contribution of our work. First, the model processes 900 uniformly sampled frames and clusters them into distinct scenes using ResNet-18, a lightweight module that measures visual correspondence between frames. Next, it extracts keywords from a given natural language query via LLM and selects the frames most relevant to those keywords using CLIP-B/16. Finally, the selected frames are described in natural language by a more powerful and computationally intensive VLM. Finally, an LLM processes the language descriptions of the selected frames to answer a given query.
@@ 3.1 Background
Recent approaches utilizing LLMs for long video question answering (LVQA) [52, 36, 16, 30, 37] can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation. Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g. frame selection in some is simply uniform sampling). In our work, we also follow this structure, but we focus on improving the frame selection stage. Under such a framework, our proposed HKS can serve as plug-in modules to replace the frame selection stage and the later two stages are similar to these prior works.
Figure 2: Detailed architecture of Hierarchical Keyframe Selector: Input frames first go through a Temporal Scene Clustering (TSC) mechanism, which divides the long-video into scenes, enabling per-scene subsampling. Next, a Coarse Keyframe Detector (CKD) is applied to select frames best-aligned with keywords relevant to the query. Finally, a Fine Keyframe detector further subsamples frames, by refining keyword alignements through a templated visual prompting.
@@ 3.2 Architecture
Consider a long video, $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$ with T, C, H, W for frames, channels, height, and width respectively. Its paired natural language query is referred as $\mathbf{q}$. Also consider a frame in $\mathbf{x}$ at timestamp $t$ as $\mathbf{x}[t]\in\mathbb{R}^{C\times H\times W}$. Our goal is to output the suitable response to this query based on information contained in the video. We refer this correct response as $\mathbf{r}$.
Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$. The HKS module initially processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$. Each of these $T^{\prime}$ frames is then passed through the captioning stage of our VLM to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$. Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$.
Our unique contributions here are the Hierarchical Keyframe Selector (HKS) and optimal prompts and description templates for the LLM when generating the final response which is also important as suggested by our experiments. The overall architecture is given in Figure 1 (left).
@@ 3.3 Hierarchical Keyframe Selector
Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema [24]. We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 390 frames. Next, the Coarse Keyframe Detector selects only 32 frames out of them, based on the alignment with keywords (Here, keywords are extracted based on answer options, via an LLM). Such coarse keyframes are then ranked based on the combination of confidence value and temporal span, and grouped into four sets, each containing eight frames. These sets are then processed through visual templating (i.e. simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
@@
We now describe our proposed Hierarchical Keyframe Selector (HKS) module. As illustrated in Figure 2, our HKS comprises of three submodules, namely 1) Temporal Scene Clustering (TSC), 2) Coarse Keyframe Detector (CKD), and 3) Fine Keyframe Detector (FKD). The inputs to HKS contain $T$ frames with each of the three stages TSC, CKD, and FKD reducing this frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
@@
Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling. The established approach for preliminary frame selection is uniform sampling. In contrast, TSC extracts per-frame visual features using a deep neural network followed by an iterative clustering procedure to identify $n$ non-overlapping frame sets. Within each of the n sets, we uniformly sample $\leq\tau$ frames obtaining a total of $T_{a}\leq\tau\times n$. Our iterative clustering procedure is outlined in Algorithm 1. It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based grouping operations.
Our reasoning for TSC is that videos, especially long-form ones, contain extensive information redundancy across its temporal dimension. For example, in a long sub-sequence of mostly static frames from a single scene, beyond its first frame the rest may contain little additional information. We hypothesize that our TSC module will utilize frame level feature similarity to cluster the frames, such that each cluster corresponds to a common scene or motion. A single cluster could contain just one frame or significantly more based on frame feature similarities. This leads to a non-uniform sampling of frames across the entire video. In contrast to uniform sampling across the entire videos where each timestamp is treated equally, in our formulation the uniform sampling within segments will allow more frames to be sampled from information heavy temporal regions.
Coarse Keyframe Detector (CKD): Unlike the TSC in prior stage, CKD reasons across both visual and language modalities (using the paired textual query) to further sub-sample the $T_{a}$ frames output from the TSC module. The inputs to this CKD stage are the $T_{a}$ frames and the paired query $\mathbf{q}$. The module generates a set of keywords $K_{\text{CKD}}$, selects $T_{b}$ frames conditioned on these keywords, and outputs the selected $T_{b}$ frames with a confidence value and keyword assigned to each frame.
@@
The CKD contains three elements: a keyword generation strategy, a dual-encoder image-text model, and an iterative algorithm for similarity based confidence assignment. The keyword generation strategy utilizes the given query, q alongside a hand-crafted templating operation or an LLM to select or generate suitable keywords. The dual-encoder image-text model uses a spatially aware contrastive langauge image pre-training (CLIP) network from [29]. For confidence assignment, we construct an algorithm as outlined in Algorithm 2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$).
The keyword generation strategy converts the query $\mathbf{q}$ into a set of keywords, $K_{\text{CKD}}$ where the number of keywords $|K_{\text{CKD}}|\in[8,25]$. The dual-encoder model converts the keyword set $K_{\text{CKD}}$ and the frame set (containing $T_{a}$ frames) into two sets of embeddings. The embeddings of each set are comparable to each other using a cosine similarity distance metric given the training formulation of the CLIP model. These two embedding sets are processed by our confidence assignment algorithm in Appendix A.2 which assigns each of the $T_{a}$ frames a confidence score and a keyword. Note that the same keyword can be assigned to multiple frames. However, given the much higher frame count compared to the keyword count and the cross-frame redundancy, we do not assign multiple keywords to a single frame.
@@
The key role of this module is language conditioned frame selection. For a single query, there can be multiple regions in a video that are highly informative but not useful or relevant in answering that query. A single query can also contain multiple different concepts and attributes that must be given attention to construct a correct answer: the keyword generation attempts to capture each of these distinct attributes. On the visual modality, a single frame will also encode multiple concepts and attributes. Our design choice for the spatially aware CLIPpy dual-encoder VLM from [29] is motivated by this nature of individual frames. Finally, confidence assignment takes into account these multiple modes of information within each frame and the query and suitably assigns a confidence score to each frame that reflects its relevance to the query. We also highlight how the confidence scores are directly linked to the related keyword (i.e. reason that makes the frame relevant), leading to better interpretability and the ability to perform further keyword-based refinement in later stages.
@@
Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame. We further illustrate this process in Figure 3 (see Visual Templating column). Each of these visual templated images also contain a subset of the keywords that correspond to their 8 images. These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix A.3 for details) are input to the VLM. The output of the VLM is used to select a subset of each 8 image group. These frames are collected as the output of the FKD stage, overall resulting in $T_{c}$ frames.
@@
3.4 Comparison with Other Keyframe Selection Methods:
We aim to highlight the main advantage of the Hierarchical Keyframe Selector over other existing keyframe selection methods. Models like VideoAgent and TraveLER provide useful comparisons, as they work with a similar-scale number of captions. VideoAgent and TraveLER rely on uniform frame selection in the first iteration without analyzing the entire video. They identify important segments based solely on these initial frames and the LLM’s response, which can be problematic if the initial uniformly selected frames are not representative of the entire video or if the LLM misinterprets the captions and prompts. In such cases, the LLM might incorrectly identify segments for further analysis. If the LLM fails to pinpoint the correct segment initially, the entire process can break down because subsequent frames will be similar to the first set, leading the LLM to continuously select frames within or near the initial segment. Additionally, for videos that are as challenging or more difficult than EgoSchema in terms of temporal complexity and activities, these models may require numerous iterations to finalize keyframes selection. This results in higher computational and latency costs, as it necessitates numerous runs of resource-intensive VLM and LLM models.
In contrast, our method analyzes the entire video with high frame rates using a lightweight ResNet-18 [12] and segments the video non-uniformly based on scene continuity. We then select several frames in each segment by measuring feature similarity between frame features and keywords using the CLIP-B/16 (0.12B) [29] which is lighted than VideoAgent’s EVA-CLIP-8Bplus (8B). By reviewing the entire video and non-uniformly selecting keyframes based on scene continuity and similarity scores, these keyframes accurately represent the question-based important frames distribution in the entire video. Furthermore, we use VLM for a fine-grained selection of keyframes, improving keyframe selection when CLIP-B/16 struggles to understand detailed atomic activities in the frames. By hierarchically segmenting the video with different modules, the resulting segments and keyframes are more reliable than those from VideoAgent. Even with more challenging videos, our process only needs to go through the video once to collect keyframes, maintaining computational efficiency.

### v5

@@ Abstract
Long-form videos that span across wide temporal intervals are highly information redundant and contain multiple distinct events or entities that are often loosely related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature explore use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector. Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, and IntentQA, while also demonstrating a strong performance on videos up to an hour long in VideoMME. The code can be found at https://github.com/jongwoopark7978/LVNet
Too Many Frames, Not All Useful:
Efficient Strategies for Long-Form Video QA
Jongwoo Park 1  Kanchana Ranasinghe 1  Kumara Kahatapitiya 1 Wonjeong Ryu2  Donghyun Kim 3  Michael S. Ryoo 1 1Stony Brook University   2KAIST AI   3Korea University jongwopark@cs.stonybrook.edu
@@ 1 Introduction
Video understanding is a long-standing vision problem (Aggarwal and Ryoo, 2011) with numerous real-world applications. It has been traditionally studied even before the era of differentiable representation learning, with hierarchical approaches focusing on longer videos (Allen and Ferguson, 1994; Ivanov and Bobick, 2000; Shi et al., 2004; Hongeng et al., 2004; Ryoo and Aggarwal, 2006). Today, video understanding research involving the language modality is particularly popular, with tasks such as video question answering that involve generating human-style conversations in response to questions regarding videos (Tapaswi et al., 2016; Zeng et al., 2017; Xu et al., 2017).
Figure 1: LVNet achieves state-of-the-art performance on EgoSchema subset while utilizing only a fraction of captioned frames. In particular, LVNet obtains its highest accuracy of 68.2% with 12 captions (VLM calls), outperforming VideoAgent and VideoTree, models using a similar-scale captions, by +8% and +5.7% (more details in Section 4.3).
@@
Recent popularity of vision-language models (VLMs), particularly approaches connecting large language models (LLMs) to vision architectures (Liu et al., 2023; Li et al., 2023b; Dai et al., 2023), has resulted in significant improvements across visual question answering (VQA) tasks. These models demonstrate exceptional performance within the image domain, and their video variants (Yu et al., 2023; Papalampidi et al., 2023; Maaz et al., 2023) perform similarly on shorter videos, yet demonstrate limited performance on long-form video benchmarks (Mangalam et al., 2023; Kahatapitiya et al., 2024; Rawal et al., 2024). This can be attributed to the nature of long-form video benchmarks, which require both temporal sequence awareness and causal reasoning. An alternate line of works (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b) adapt LLMs that contain strong reasoning abilities for this task, using image VLMs to generate per-frame natural language descriptions, followed by video question answering purely within the language domain. However, these methods employ expensive VLMs to caption a large number of uniformly sampled frames. Such a design choice leading to high compute expense, is questioned in (Buch et al., 2022; Ranasinghe et al., 2024; Wang et al., 2024b), and is the key motivation for our exploration of key frame selection, i.e. identifying a minimal set of frames most useful for correctly answering a given video-question pair.
@@
Therein, we propose LVNet, a framework containing a novel Hierarchical Keyframe Selector (HKS) that performs efficient key-frame selection followed by VLM and LLM for caption and answer generation as illustrated in Fig. 2. Aligned with prior work (Zhang et al., 2023; Wang et al., 2024d, b), the per-frame captions are processed with a powerful LLM to generate correct answers for a given video-question pair. As shown in Fig. 1, LVNet achieves strong performance using a small set of keyframes from the HKS. The scope of this work focuses on optimizing the prior two stages. We summarize our key contributions as follows:
1. Hierarchical Keyframe Selector (HKS): The proposed HKS consists of three submodules for efficient keyframe selection.
  1. (a)
    Temporal Scene Clustering (TSC)
    • Performs non-uniform frame sampling by clustering visually similar frames.
    • Reduces redundancy in long videos while capturing key scenes.
    • A lightweight module for efficient filtering of dense frames.
  2. (b)
    Coarse Keyframe Detector (CKD)
    • Generates keywords representing atomic activities using the given query and an LLM.
    • Assigns confidence scores to frames based on keyword relevance.
    • Samples high-confidence frames for improved interpretability over visual-only selection.
  3. (c)
    Fine Keyframe Detector (FKD)
    • Refines frame selection by combining multiple frames using visual templating and a VLM.
    • Enables higher-level reasoning and natural language-based selection.
    • Achieves better accuracy than CKD’s keyword-based selection.
@@
2. Zero-Shot Long-Form Video Understanding: Our framework operates zero-shot without requiring video-level training. This makes it highly efficient for long-form video understanding.
Figure 2: (top) Overview: LVNet uses a Hierarchical Keyframe Selector (HKS) module to select keyframes, followed by VLM & LLM for caption and answer generation. (below) HKS Module processes dense frames with lighter modules and progressively exploits heavier, more performance-oriented modules on smaller subsets of frames to ensure efficient computation.
@@
Proposed LVNet achieves state-of-the-art results compared to models utilizing similar number of captions on three long-form video question answering benchmarks— EgoSchema, NExT-QA, and IntentQA(Sec. 4.2). This demonstrates strong performance and generality of our approach.
@@ 2 Related Work
Video Question Answering: Visual question answering (VQA) involves generating open-ended textual content conditioned on an image and natural language query (Agrawal et al., 2015). Its video variant, Video-VQA (Yu et al., 2019a) replaces images with videos. Multiple early datasets focus on querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019a). Later tasks require explicit temporal understanding of sequential events (Lei et al., 2018, 2020; Yu et al., 2019b). More recent datasets focus on longer videos containing multiple actions and scenes spread over wide time intervals (termed long-form videos) (Xiao et al., 2021; Li et al., 2022). Referred to as long-form video question answering (LVQA), these benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023). Some works tackling such video VQA tasks leverage graph networks to model cross object / event relations (Hosseini et al., 2022; Xiao et al., 2022a, b). A more recent line of works integrate LLMs to tackle this task (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Wang et al., 2024b; Ranasinghe et al., 2024; Wang et al., 2024d; Fan et al., 2024) utilizing the strong reasoning skills of LLMs. A common aspect is the use of a vision language model (VLM) to convert frame level visual information into natural language. This in turn is input to the LLM which makes a final prediction.
Unlike these methods, LVNet incorporates a unique Hierarchical Keyframe Selector that progressively reduces the number of keyframe candidates. Lighter modules are applied to dense frames, while heavier, more performance-focused modules are applied to a small subset of filtered frames. Additionally, LVNet does not require video-level training, unlike earlier supervised approaches.
Feature Ours VA Tr. VT
(effective selection)
Uses non-uniform sampling
Scene continuity-based selection
Robust to initial frames
Fine-grained visual refinement
(compute efficient)
Lightweight feature extraction
Single pass inference
Table 1: LVNet exhibits unique features compared to VideoAgent (VA) Fan et al. (2024), Traveler (Tr.) Shang et al. (2024) and VideoTree (VT) Wang et al. (2024e). See Appendix A.5 for details.
Frame Selection in Videos: The task of frame selection in videos has been long explored in video (Davis and Bobick, 1997; Zhao et al., 2017) with more recent works focused directly on long-form video question answering (Buch et al., 2022; Wang et al., 2024e; Fan et al., 2024). Most similar to our work is (Wang et al., 2024b) which employs an LLM based strategy for video frame selection. However, our LVNet differs with several unique features as summarized in Table 1.
@@ 3 Method
Our proposed LVNet comprises of three components: a Hierarchical Keyframe Selector (HKS), a Vision Language Model (VLM), and a Large Language Model (LLM) as illustrated in Figure 2. The HKS, an efficient, hierarchical keyframe selector, is the core contribution of our work. First, the model processes 900 uniformly sampled frames and clusters them into distinct scenes Next, it extracts keywords from a given natural language query via LLM and selects the frames most relevant to those keywords. Finally, the selected frames are described in natural language by a more powerful and computationally intensive VLM. Finally, an LLM processes the language descriptions of the selected frames to answer a given query.
@@ 3.1 Background
Recent approaches utilizing LLMs for long video question answering (LVQA) (Zhang et al., 2023; Wang et al., 2023; Kahatapitiya et al., 2024; Ranasinghe et al., 2024; Wang et al., 2024b) can be viewed as a composition of three sequential stages: a) frame selection, b) VLM based frame captioning, and c) LLM based answer generation. Note that the complexity of each stage varies across methods given their focus on different aspects of the LVQA task (e.g. frame selection in some is simply uniform sampling). In our work, we also follow this structure, but we focus on improving the frame selection stage. Under such a framework, our proposed HKS can serve as plug-in modules to replace the frame selection stage and the later two stages are similar to these prior works.
Figure 3: Qualitative example: We illustrate a challenging long-video QA scenario from EgoSchema (Mangalam et al., 2023). We consider an input of 900 frames, which first get clustered into scenes and subsampled to retain around 390 frames. Next, the Coarse Keyframe Detector selects only 32 frames out of them, based on the alignment with keywords (Here, keywords are extracted based on answer options, via an LLM). Such coarse keyframes are then ranked based on the combination of confidence value and temporal span, and grouped into four sets, each containing eight frames. These sets are then processed through visual templating (i.e. simple concatenation across space) and fed into a VLM for Fine Keyframe Detection, resulting in just 12 frames.
@@ 3.2 Architecture
Consider a video, $\mathbf{x}\in\mathbb{R}^{T\times C\times H\times W}$ with T, C, H, W for frames, channels, height, width respectively and its paired natural language query $\mathbf{q}$. Also consider a frame in $\mathbf{x}$ at timestamp $t$ as $\mathbf{x}[t]\in\mathbb{R}^{C\times H\times W}$. Our goal is to output a response, referred as $\mathbf{r}$, suitable for the given query $\mathbf{q}$ based on information contained in the video $\mathbf{x}$.
Our LVNet processes a given video-query ($\mathbf{x},\mathbf{q}$) pair to output a response, $\mathbf{\hat{r}}$. The HKS module initially processes this video-query pair, selects $T^{\prime}$ keyframes, and outputs a deterministically sub-sampled video $\mathbf{x^{\prime}}\in\mathbb{R}^{T^{\prime}\times C\times H\times W}$. Each of these $T^{\prime}$ frames is then passed through the captioning stage of our VLM to generate a set of natural language descriptions, $D=\{d_{1},d_{2},...d_{T^{\prime}}\}$ where $d_{i}$ describes the frame $\mathbf{x^{\prime}}[i]$. Finally, the LLM processes all descriptions $D$ and the query $\mathbf{q}$ to generate response $\mathbf{\hat{r}}$. We illustrate this overall architecture in Figure 2.
@@ 3.3 Hierarchical Keyframe Selector
We now describe our proposed Hierarchical Keyframe Selector (HKS) module. As illustrated in Figure 2, our proposed HKS comprises of three sequential submodules, each reducing the frame count to $T_{a}$, $T_{b}$, and $T_{c}=T^{\prime}$ respectively.
@@
Temporal Scene Clustering (TSC): The role of TSC is to perform visual content aware preliminary frame sampling. The established approach for preliminary frame selection is uniform sampling (limited to at most 200 frames). In contrast, TSC processes 900 to 1800 uniformly sampled frames to extract per-frame visual features using a lightweight deep neural network (ResNet-18) followed by a clustering procedure to identify $n$ non-overlapping frame sets. Within each of the n sets, we uniformly sample $\leq\tau$ frames obtaining a total of $T_{a}\leq\tau\times n$. Our iterative clustering procedure is outlined in Algorithm 1. It calculates pairwise distances between all frames accounting for intra-frame local information using the extracted per-frame features, followed by n iterative frame similarity based clustering operations. A single cluster could contain just one frame or significantly more based on frame feature similarities, leading to a non-uniform sampling of frames across the entire video. This allows more frames to be sampled from the information heavy temporal regions of videos.
@@
Coarse Keyframe Detector (CKD): Unlike TSC in the prior stage, CKD reasons across both visual and language modalities (using the paired textual query, $q$) to further sub-sample $T_{a}$ into $T_{b}$ frames. CKD contains three elements: keyword generation strategy, dual-encoder image-text model, and similarity based confidence assignment algorithm. Keyword generation utilizes the given query, q, alongside hand-crafted templating operations or an LLM to select or generate suitable keywords. The dual-encoder image-text model uses a spatially aware contrastive language image pre-training (CLIP) network from (Ranasinghe et al., 2023). For confidence assignment, we construct an algorithm as outlined in Algorithm 2 which processes two lists, one of frames and one of keywords, and then calculates their pairwise likelihood of occurrence to assign each frame a confidence value (that reflects its usefulness to answer the query, $\mathbf{q}$). See  Appendix A.3 for more details.
@@
For a single query, there can be multiple regions in a video that are highly informative but not useful or relevant in answering that query. A single query can also contain multiple different concepts and attributes that must be given attention to construct a correct answer: the keyword generation attempts to capture each of these distinct attributes. On the visual modality, a single frame will also encode multiple concepts and attributes. Our design choice for the spatially aware CLIPpy dual-encoder VLM from (Ranasinghe et al., 2023) is motivated by this nature of individual frames. Finally, confidence assignment takes into account these multiple modes of information within each frame and the query to suitably assign confidence scores to each frame that reflects its query relevance. We also highlight how the confidence scores are directly linked to the related keyword (i.e. reason that makes the frame relevant), leading to better interpretability and the ability to perform further keyword-based refinement in later stages.
@@
Our visual templating module partitions the $T_{b}$ frames into sets of 8 ordered by their confidence scores, arranges frame sets as grids to form a collage-style image, and annotates that image with visually identifiable tags corresponding to each frame. We further illustrate this process in Figure 3 (see Visual Templating column). Each of these visual templated images also contain a subset of keywords that correspond to their 8 images. These resulting visual templated images along with a prompt containing their associated keywords and instructions to select a frame subset based on valid association between keywords and images (see Appendix A.4 for details) are input to the VLM. The output of the VLM is used to select a subset of each 8 image group. These frames are collected as output of the FKD stage, overall resulting in $T_{c}$ frames.